Jasmine spends her working hours (8 hours) filing, typing, and receiving calls. If she approximately

performs these functions in the ratio 3:4:5, how many hours does she spend for each function

respectively?


1
Expert's answer
2021-10-11T09:48:35-0400
"3+4+5=12(parts)"

filing:

"\dfrac{8\ h}{12}\cdot 3=2\ h"

typing

"\dfrac{8\ h}{12}\cdot 4=2\dfrac{2}{3}\ h=2\ h\ 40\ min"

receiving calls


"\dfrac{8\ h}{12}\cdot 5=3\dfrac{1}{3}\ h=3\ h\ 20\ min"